OpenStudy (anonymous):

How do I calculate arc length? Imagine a slice of pizza, I need to calculate the area of the slice, the interior angle, and the arc length of the slice. How would I start it? (I have to repeat this 7 more times, just want to know how to get started)

OpenStudy (anonymous):

It's actually really easy! The formula for arclength if you have the function is: \[\int\limits_{}^{} \sqrt{1+\frac{ dy }{ dx }}dx\] or if you have the parametric equation then it is: \[\int\limits\limits_{}^{} \sqrt{\frac{ dx ^{2} }{ dt }+\frac{ dy ^{2} }{ dt }}dt\]

OpenStudy (anonymous):

Thank you! :D If I could though, I need a bit more help deciphering the formula you gave me. What does dx, dt, and the symbol next to the square root sign mean?

OpenStudy (anonymous):

Take the integral of the squareroot with respect to x or respect to t with the second formula

OpenStudy (anonymous):

Integral?

OpenStudy (anonymous):

This is highschool Geometry by the way.

OpenStudy (anonymous):

This is calculus 3? Multivariable calculus?

OpenStudy (anonymous):

My vocabulary may be big but it isn't THAT big haha.

OpenStudy (anonymous):

Calculus 3? I had no idea that even existed xD. I doing... ehh "somewhat" simple Geometry

OpenStudy (anonymous):

I hope I'm not wasting your time by the way x'c

OpenStudy (anonymous):

Ahh! Ok I thought this was calc 3! The equation for arc length if you want to do it the basic way is: (2pi*r)*(n degrees/360) But if you want to impress your teacher you could do some multivariable calculus to solve the equation!

OpenStudy (anonymous):

Okay so I have the formula written down for arc length. For the interior angles though, I've always been rusty with that kind of stuff. Would I determine the interior angles by the arc length?

OpenStudy (anonymous):

It makes sense if that's how

OpenStudy (anonymous):

If you have the radius and arc length then you could find out the inner angle!

OpenStudy (anonymous):

:D yay! Well check it out: (bare with me, I gotta type it lol)

OpenStudy (anonymous):

I have a pizza that is 20 inches in diameter. I'm going to cut it into 8 slices, for the first step I have to find the interior angles and arc lengths of EACH slice. Since they are going to be congruent I only have to do it for ONE, I already have the tools to do that (My many thanks). I'm going to plug in what you told me and solve the first step. How long are you going to be on?

OpenStudy (anonymous):

also, "n" is that number of degrees or angles amongst the 8 slices?

OpenStudy (anonymous):

We know that a circle is 360 degrees and there are 8 slices so each slice is (8/360) degrees! The radius is 10 because 20/2 is 10. Just plug into the formula to get the answer!

OpenStudy (anonymous):

Okie dokie, stand by.

OpenStudy (anonymous):

I got 1.395 repeating

OpenStudy (anonymous):

So my arc length of EACH slice will be 1.4? or 2?

OpenStudy (anonymous):

Yep! It would be 1.396 or (4/9)pi

OpenStudy (anonymous):

yesss! okay now I gotta do the AREA of the slice

OpenStudy (anonymous):

I imagine that's going to have something to do with one-eighth and Cirumference, along with pi haha

OpenStudy (anonymous):

The area is different than circumference! The area is Pi r^2, but the circumference is the derivative: 2Pi*r

OpenStudy (anonymous):

Hey I have to go eat dinner, I am extremely grateful your helping me but I haven't seen my dad in 4 months so would you be willing to resume this innnnn like 15-20 minutes? Your working magic right now and I still need your help :P lol

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

:D yay! Alright, I'll be back in 15!

OpenStudy (anonymous):

Aye you there?

OpenStudy (anonymous):

Love it lol, reason for it?

OpenStudy (anonymous):

no reason. What's the problem?

OpenStudy (anonymous):

lol nothing haha.

OpenStudy (anonymous):

my dad said that instead of 8/360 it would be 360/8. for the arc formula

OpenStudy (anonymous):

? Hello

OpenStudy (anonymous):

You divide the total degrees of the circle (360) by the number of pieces which is 8. So yep he's right I wrote it out wrong. Once you have that, then you plug it into n

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

Lemme redo it then.

OpenStudy (anonymous):

Sorry about that!

OpenStudy (anonymous):

It's all good! :D hold on

OpenStudy (anonymous):

I got 2826

OpenStudy (anonymous):

It would be 2.5 Pi or 7.853

OpenStudy (anonymous):

First d 360/8 to get the inner angle, then plug that into n and solve!

OpenStudy (anonymous):

so 360/8= n?

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

okay hold on

OpenStudy (anonymous):

okay so now i have 62.8 x 360/45. correct?

OpenStudy (anonymous):

What??? Where did these numbers come from?! Most teachers would rather have you write the symbol for pi instead of multiplying it out

OpenStudy (anonymous):

D: im so confused doe. I need verbal exchange for this. I'm sorry :/ your help has been appreciated though! You get medals lol. Ill jsut have my dad help me. Im tired anyways lol

OpenStudy (anonymous):

Hope all works out!

OpenStudy (anonymous):

Thank you! :D